OpenStudy (anonymous):

Solve the compound inequality 7>-4x+6 or 12≤-3x+5

OpenStudy (anonymous):

solve the two inequalities separately \[7>-4x+6\] \[4x>6-7\] \[4x>-1\] \[x>-\frac{1}{4}\]

OpenStudy (saifoo.khan):

.

OpenStudy (anonymous):

or \[12\geq -3x+5\] \[7\geq -3x\] \[3x\geq -7\] \[x\geq -\frac{7}{3}\]

OpenStudy (anonymous):

now since your question asks for OR, you have \[x>-\frac{1}{4}\] OR \[x\geq -\frac{7}{4}\] so take the biggest set possible which means \ \[x\geq - \frac{7}{4}\]

OpenStudy (anonymous):

so i have a question there are some options does the U mean or?